OpenStudy (anonymous):

Which of the binomials below is a factor of this trinomial? -3x2 - 21x + 24

OpenStudy (gorv):

-3(x2+7x-8)=-3(x2+8x-x-8) =-3(x(x+8)-1(x+8)) =-3(x+8)(x-1)